Directions: find a common solution for each system of equations. The equations are y=x+1 and y=3x-2. My answer for the equations was (1.5,2.5). Am I correct. If not show how you got the answer. Thanx.

Answer by checkley71(8403)

X+1=3X-2 WE CAN SET THESE 2 EQUATION = BECAUSE THEY BOTH ARE EQUAL TO Y.
X-3X=-2-1
-2X=-3
X=-3/-2
X=3/2 OR 1 1/2 OR 1.5 ANSWERS.
Y=1.5+1
Y=2.5 ANSWER.
PROOF.
Y=3*1.5-2
Y=4.5-2
Y=2.5 ANSWER.